OpenStudy (anonymous):

Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution?

OpenStudy (anonymous):

I will show you

OpenStudy (anonymous):

okay thankyou

OpenStudy (anonymous):

so, i will give you an equation that i created okay

OpenStudy (anonymous):

i will show you the process on how to solve it

OpenStudy (anonymous):

but, you should not just copy my work, try and think of your own

OpenStudy (anonymous):

Consider the equation sqrt(-2x + 3) = x To solve this, first we would square both sides. -2x + 3 = x^2. Getting everything to one side, we are trying to solve x^2 + 2x - 3 = 0. Factoring, we get (x - 1)(x + 3) = 0. Setting each term equal to 0, we see that we get possible solutions x = 1 and x = -3. Finally, we plug back in to the original equation to see if these are true solutions. When x = 1, we get sqrt(-2 + 3) = sqrt(1) = 1, which means 1 is a solution. When x = -3, we get sqrt(-2(-3) + 3) = sqrt(9) = 3, which is NOT -3, so -3 is not a solution. The only solution to this radical equation is x = 1.

OpenStudy (anonymous):

okay thanks.